# algorithmic

### v1

@@ 8.1.4 SEAL-type Methods
SEAL-type methods have shown superior performance among existing link prediction approaches [21, 36, 36]. SEAL and its subsequent works [74, 37, 60, 70, 59] address the link prediction problem by classifying the subgraphs that are extracted specifically for candidate links. SEAL [74] extracts a local enclosing subgraph for each candidate link and uses a GNN [75] to classify these subgraphs for link prediction. GraiL [60] is developed for inductive link prediction. It is similar to SEAL but it replaces SortPooling [75] with MEAN-pooling. DEGNN [37] proposes a distance encoding GNN. Cai et al. [6] transform the enclosing subgraph into a corresponding line graph and address the link prediction task with the node classification problem in its line graph. Pan et al.[46] follow the subgraph strategy in SEAL while designing a new pooling mechanism called WalkPool. SUREL [70] proposes an algorithmic technique to improve the computational efficiency of subgraph generation in SEAL. SIEG [3] incorporates the structural information learned from the enclosing subgraphs into the GNN for link prediction, which fuses topological structures and node features to take full advantage of graph information for link prediction.
8.1.5 Methods Specific for Link Prediction
@@
Various link prediction-specific methods have been introduced [77, 27]. Wang et al. [63] present PLNLP by jointly using the representations learned by a GNN, distance encoding, etc. Neo-GNN [73] weighted aggregates the link prediction scores obtained by heuristics and a GNN. NBFNet [82] generalizes traditional path-based link heuristics into a path formulation. Singh et al. [56] show that adding a set of edges to the graph as a pre-processing step can improve the performance of link prediction models. PermGNN [50] optimizes the neighborhood aggregator directly by link samples. Zhao et al. [78] study counterfactual questions about link existence by causal inference. RelpNet [64] aggregates edge features along the structural interactions between two target nodes. Guo et al. [18] propose cross-model distillation techniques for link prediction. Shang et al. [55] propose a negative link sampling method PbTRM based on a policy-based training method. Li et al. [35] study the integration of large language models (LLMs) and prompt learning techniques with graphs, enhancing graph transfer capabilities across diverse tasks and domains.
8.2 Limitation Analysis
@@
We provide two basic insights into the application of GNNs in link prediction. Firstly, we show that aggregation-based GNNs inherently lack the ability to learn NCN-dependent structural information for link prediction (Remark 3). Secondly, we demonstrate that node embeddings can boost the performance of GNN-based link prediction models on dense graphs (Remark 4 and Finding 5). These can serve as effective avenues to identify and interpret the limitations of existing link prediction methods. To illustrate this, we present two case studies.
@@
Figure 3: The algorithm flow of SEAL-type link prediction methods.
@@
Case study 1. Can SEAL effectively learn NCN-dependent structural information? SEAL-type methods have achieved the best performance on several link prediction datasets [74, 37, 60, 70]. SEAL [74] has proven that most link heuristics between two nodes can be computed approximately within an enclosing subgraph extracted specifically for that two nodes. As shown in Fig. 3, most SEAL-type methods employ GNNs for graph representation learning, with the expectation that from such enclosing subgraphs, the GNN can learn the structural information equivalent to link heuristics including CN, AA, Katz, etc. However, whether this expectation holds true has not been thoroughly investigated in existing works. Herein we present a rough analysis to examine this issue.
@@
First, the GNNs used in SEAL-type methods, e.g., DGCNN [75] in SEAL [74], R-GCN [52] in GraiL [60], still belong to the type of aggregation-based GNNs. According to Remark 3, these GNNs inherently cannot effectively learn NCN-dependent structural information.
Furthermore, we have noticed that SEAL-type methods typically use a labeling technique [76] to add labeling features to each node in the enclosing subgraph. The labeling features of each node describe the relationship of that node to the target two nodes. Fig. 4 illustrates such a labeling method, where the labeling features of a node are the shortest path distances from the node to the target pair of nodes. The work [76] points out that the labeling features can help the GNN learn the structural information about the number of common neighbors. Their explanation is as follows. As shown in Fig. 4, for node $v$ and $u$, in the first iteration of the neighborhood aggregation of a GNN, only the common neighbors between node $v$ and $u$ will receive the labeling messages from both $v$ and $u$; then in the second iteration, the common neighbors will pass such messages back to both $v$ and $u$, which can encode the number of common neighbors into the representations of node $v$, $u$.
However, the aforementioned explanation raises questions regarding its validity. In the second iteration, apart from the common neighbors, the non-common neighbors of node $v$ also pass their messages back to $v$. The messages from all neighbors of $v$ are then aggregated through a set-based pooling (e.g., MEAN or attention-based pooling as shown in Eq. 2). Such an aggregated result for node $v$ would wash out the distinguishable labels’ information. We present an example to illustrate this. As shown in Fig. 4, if the pooling method in a GNN is MEAN, then the aggregation of the labeling features of the neighbors of node $v$ would be equal to that of node $w$, i.e., $\mathrm{MEAN}(\{10,3,2,1\})=\mathrm{MEAN}(\{4,6,4,3,3\})$. This means that the distinct labeling features of the neighbors of a node are not effectively kept in the aggregated result. In other words, the aggregated results for node $v$ and $w$ in the positive and negative link samples become indistinguishable. Note that attention-based pooling in GNNs like GAT [61] also suffers from the above limitation for the same reason as the analysis of Remark 2. The same goes for node $u$. It should be noted that our example is merely for illustrative purposes. In practice, a GNN layer contains a series of complicated operations such as linear and non-linear transformations, dropout, residual connection, and others. The structural information in the labeling features could be partially kept in the learned representations.
@@
Figure 4: Node labeling in SEAL-type methods. The left is a subgraph specific for a positive link sample and the right is a negative one. The labeling features are based on the SPDs from every node (here only show the first-order neighbors of node $v$ or $w$) to the target pair of nodes. For example, on the left, the node with the labeling $(1,10)$ indicates that the SPD from this node to node $v$ and $u$ is $1$ and $10$, respectively.
@@
Table 4: Results on OGB link prediction datasets. Higher scores indicate better performance, with the best results highlighted in bold. Herein, "HE" stands for Heuristic Encoding, "X" represents node attributes, and "NE" denotes Node Embedding.
@@
MF [44] $13.68\pm 4.75$ $38.86\pm 0.49$ $32.29\pm 0.94$ $51.86\pm 8.43$
FSSDNMF [8] $14.62\pm 2.64$ $37.95\pm 3.25$ $34.15\pm 1.16$ $54.71\pm 8.73$
DeepWalk [47] $26.42\pm 6.10$ $50.37\pm 0.34$ $28.88\pm 1.63$ $60.11\pm 0.23$
Node2vec [17] $23.26\pm 2.09$ $48.88\pm 0.54$ $22.26\pm 0.83$ $61.41\pm 0.11$
NodePiece  [14] $24.15\pm 3.04$ $47.88\pm 0.41$ $22.85\pm 0.94$ $61.52\pm 2.91$
GraphSAGE [19] $83.90\pm 4.74$ $48.10\pm 0.81$ $16.55\pm 2.40$ $82.60\pm 0.36$
GAT [61] $95.38\pm 0.94$ $52.26\pm 0.85$ $51.33\pm 2.16$ $83.17\pm 0.54$
Neo-GNN [73] $75.72\pm 3.42$ $55.31\pm 0.53$ $49.13\pm 0.60$ $87.26\pm 1.84$
PLNLP [63] $90.88\pm 3.13$ $52.92\pm 0.98$ $32.38\pm 2.58$ $84.92\pm 0.29$
NBFNet [82] $18.14\pm 2.12$ $51.15\pm 1.38$ $23.96\pm 2.03$ $74.91\pm 2.37$
SEAL [74] $30.56\pm 3.86$ $54.71\pm 0.79$ $48.80\pm 4.56$ $87.67\pm 0.32$
DEGNN [37] $26.63\pm 6.42$ $53.74\pm 0.45$ $36.48\pm 5.38$ $60.30\pm 0.81$
SIEG [3] $31.95\pm 3.93$ $55.35\pm 0.52$ $53.35\pm 1.39$ $\mathbf{89.87\pm 0.10}$
@@
Case study 2. NBFNet lacks the algorithmic ability to leverage node embeddings. NBFNet [82] is a model specifically developed for link prediction. Differently from GNN-based link prediction methods like Algorithm 2 and SEAL-type methods, NBFNet generalizes traditional link heuristics such as Katz index [26], Personalized PageRank [33] into a general formulation and approximates such formulation using a special network. Unlike aggregation-based GNNs that propagate and aggregate node-wise representations, NBFNet is designed to train edge-wise representations. The model architecture of NBFNet makes it hardly consider node-wise information. This would make NBFNet lack the algorithmic ability to train powerful node embeddings and may lead to non-competitive link prediction performance on dense graphs, considering the strong performance of the GNN only using node embeddings on dense graphs as shown in Fig. 2.
8.3 Further analysis of experimental results
@@
We expand our limitation analysis of existing link prediction methods by examining their experimental performance on four OGB benchmark datasets. The results are presented in Table 4 and Fig. 5. In the interest of brevity, our analysis focuses on several main types of methods.
@@
First, as shown in Fig. 5, the performance of heuristic methods is not stable across the four datasets. For example, RA performs best on ogbl-ppa but second worst on ogbl-ddi. These results are consistent with the research of [31] which indicates that many link heuristics are designed for specific applications and may perform well only on those specific graphs. Moreover, the unstable performance of every single heuristic confirms the need of combining multiple heuristics in link prediction, as shown in our Algorithm 2.
@@
We also report the heuristic encoding results (Fig. 6) obtained through Algorithm 2 when only employing heuristic encoding. We find that it is not always true that the more the heuristics used, the better the performance of HE. In contrast, encoding a certain number of heuristics can yield the best performance, whereas encoding too many heuristics would pose an optimization challenge.
@@
The node embedding methods based on relative distance encoding (DeepWalk [47], NodePiece [14]) perform slightly better than those based on matrix factorization (MF [44], FSSDNMF [8]). Nevertheless, all these methods fall short compared to other methods. This could be attributed to the limitations of such methods, e.g., reliance solely on the adjacency matrix or unsupervised learning without link samples. This also underscores the critical role of link samples in supervising the training of node embeddings for link prediction.
@@
Fig. 5 also presents the results of MLP and general GNNs (GCN [29], GAT [61] and JKNet [68]) that use node embeddings only. MLP(NE) performs much worse than GNNs, demonstrating the significance of neighborhood aggregation of GNNs in training node embeddings, considering that MLP updates each node’s representation independently of other nodes. Furthermore, GCN(NE) and GAT(NE) perform comparably, indicating that the expressiveness of GCN is sufficient for learning node embedding. The similar performance of GCN and GAT empirically supports our analyses in Remark 2 and 3, where we point out that the attention mechanism (e.g., GAT) cannot address the inherent issue of GNNs in learning structural information related to the number of each node’s neighbors and of common neighbors between two nodes.
@@
In Fig. 5, SEAL-type methods show state-of-the-art performance. Especially, SIEG [3] achieves the best results on two sparse graphs, i.e., ogbl-collab and ogbl-citation2 with graph degrees of $8$ and $21$, respectively. However, they perform worse than general GNNs with node embeddings (GCN(NE) [29], GAT(NE) [61] and JKNet(NE) [68]) on two dense graphs, i.e., ogbl-ppa and ogbl-ddi. This discrepancy in the performance of SEAL-type methods could be attributed to the algorithmic challenge of training node embeddings using subgraphs. Unlike general GNNs, the algorithm of SEAL-type methods limits each node to perceive other nodes within the subgraph rather than the entire graph, thereby restricting the information flow between nodes and potentially reducing the efficiency of learning node embeddings.
@@
Besides, Fig. 5 shows two link prediction-specific methods, namely NBFNet [82] and Neo-GNN [73]. NBFNet underperforms on four datasets, which aligns with the limitations identified in Section 8.2. Neo-GNN predicts link likelihood by combining the scores obtained by heuristic methods and the result produced by a GNN. It performs on par with the state-of-the-art SEAL-type methods on two sparse graphs (ogbl-collab and ogbl-citation2).
@@
Lastly, our GNN(X)+HE based on Algorithm 2 performs better than SEAL-type methods on ogbl-collab, supporting our limitation analysis of SEAL-type methods in Section 8.2, i.e., such methods could not effectively learn the information equivalent to NCN-dependent heuristics.
@@
9 Implication for Practical Applications

### v2

@@ VIII-A4 SEAL-type Methods
SEAL-type methods have shown superior performance among existing link prediction approaches [hu2020open, li2024evaluating, li2024evaluating]. SEAL and its subsequent works [zhang2018link, li2020distance, teru2020inductive, yin2022algorithm, tan2023bring] address the link prediction problem by classifying the subgraphs that are extracted specifically for candidate links. SEAL [zhang2018link] extracts a local enclosing subgraph for each candidate link and uses a GNN [zhang2018end] to classify these subgraphs for link prediction. GraiL [teru2020inductive] is developed for inductive link prediction. It is similar to SEAL but it replaces SortPooling [zhang2018end] with MEAN-pooling. DEGNN [li2020distance] proposes a distance encoding GNN. Cai et al. [cai2021line] transform the enclosing subgraph into a corresponding line graph and address the link prediction task with the node classification problem in its line graph. Pan et al.[pan2021neural] follow the subgraph strategy in SEAL while designing a new pooling mechanism called WalkPool. SUREL [yin2022algorithm] proposes an algorithmic technique to improve the computational efficiency of subgraph generation in SEAL. SIEG [ai2022structure] incorporates the structural information learned from the enclosing subgraphs into the GNN for link prediction, which fuses topological structures and node features to take full advantage of graph information for link prediction.
VIII-A5 Methods Specific for Link Prediction
@@
Various link prediction-specific methods have been introduced [zhang2023page, khemani2024review]. Wang et al. [wang2021pairwise] present PLNLP by jointly using the representations learned by a GNN, distance encoding, etc. Neo-GNN [yun2021neo] weighted aggregates the link prediction scores obtained by heuristics and a GNN. NBFNet [zhu2021neural] generalizes traditional path-based link heuristics into a path formulation. Singh et al. [singh2021edge] show that adding a set of edges to the graph as a pre-processing step can improve the performance of link prediction models. PermGNN [roy2021adversarial] optimizes the neighborhood aggregator directly by link samples. Zhao et al. [zhao2022learning] study counterfactual questions about link existence by causal inference. RelpNet [wu2022relpnet] aggregates edge features along the structural interactions between two target nodes. Guo et al. [guo2022linkless] propose cross-model distillation techniques for link prediction. Shang et al. [shang2022improving] propose a negative link sampling method PbTRM based on a policy-based training method. Li et al. [li2024graph] study the integration of large language models (LLMs) and prompt learning techniques with graphs, enhancing graph transfer capabilities across diverse tasks and domains.
VIII-B Limitation Analysis
@@
We provide two basic insights into the application of GNNs in link prediction. Firstly, we show that aggregation-based GNNs lack the ability to learn NCN-dependent structural information for link prediction (Remark 3). Secondly, we demonstrate that node embeddings can boost the performance of GNN-based link prediction models on dense graphs (Remark 4 and Finding 5). These can serve as effective avenues to identify and interpret the limitations of existing link prediction methods. To illustrate this, we present two case studies.
@@
Figure 3: The algorithmic flow of SEAL-type link prediction methods.
@@
Case study 1. Can SEAL effectively learn NCN-dependent structural information? SEAL-type methods have achieved the best performance on several link prediction datasets [zhang2018link, li2020distance, teru2020inductive, yin2022algorithm]. SEAL [zhang2018link] has proven that most link heuristics between two nodes can be computed approximately within an enclosing subgraph extracted specifically for that two nodes. As shown in Fig. 3, most SEAL-type methods employ GNNs for graph representation learning, with the expectation that from such enclosing subgraphs, the GNN can learn the structural information equivalent to link heuristics including CN, AA, Katz, etc. However, whether this expectation holds true has not been thoroughly investigated in existing works. Herein we present a rough analysis to examine this issue.
@@
First, the GNNs used in SEAL-type methods, e.g., DGCNN [zhang2018end] in SEAL [zhang2018link], R-GCN [schlichtkrull2018modeling] in GraiL [teru2020inductive], still belong to the type of aggregation-based GNNs. According to Remark 3, these GNNs cannot effectively learn NCN-dependent structural information.
@@
Figure 4: Node labeling in SEAL-type methods. The left is a subgraph specific for a positive link sample and the right is a negative one. The labeling features are based on the SPDs from every node (here only show the first-order neighbors of node $v$ or $w$) to the target pair of nodes. For example, on the left, the node with the labeling $(1,10)$ indicates that the SPD from this node to node $v$ and $u$ is 11 and $10$, respectively.
Furthermore, we have noticed that SEAL-type methods typically use a labeling technique [zhang2021labeling] to add labeling features to each node in the enclosing subgraph. The labeling features of each node describe the relationship of that node to the target two nodes. Fig. 4 illustrates such a labeling method, where the labeling features of a node are the shortest path distances from the node to the target pair of nodes. The work [zhang2021labeling] points out that the labeling features can help the GNN learn the structural information about the number of common neighbors. Their explanation is as follows. As shown in Fig. 4, for node $v$ and $u$, in the first iteration of the neighborhood aggregation of a GNN, only the common neighbors between node $v$ and $u$ will receive the labeling messages from both $v$ and $u$; then in the second iteration, the common neighbors will pass such messages back to both $v$ and $u$, which can encode the number of common neighbors into the representations of node $v$, $u$.
However, the aforementioned explanation raises questions regarding its validity. In the second iteration, apart from the common neighbors, the non-common neighbors of node $v$ also pass their messages back to $v$. The messages from all neighbors of $v$ are then aggregated through a set-based pooling (e.g., MEAN or attention-based pooling as shown in Eq. 2). Such an aggregated result for node $v$ would wash out the distinguishable labels’ information. We present an example to illustrate this. As shown in Fig. 4, if the pooling method in a GNN is MEAN, then the aggregation of the labeling features of the neighbors of node $v$ would be equal to that of node $w$, i.e., $\mathrm{MEAN}(\{10,3,2,1\})=\mathrm{MEAN}(\{4,6,4,3,3\})$. This means that the distinct labeling features of the neighbors of a node are not effectively kept in the aggregated result. In other words, the aggregated results for node $v$ and $w$ in the positive and negative link samples become indistinguishable. Note that attention-based pooling in GNNs like GAT [gat2018graph] also suffers from the above limitation for the same reason as the analysis of Remark 2. The same goes for node $u$. It should be noted that our example is merely for illustrative purposes. In practice, a GNN layer contains a series of complicated operations such as linear and non-linear transformations, dropout, residual connection, and others. The structural information in the labeling features could be partially kept in the learned representations.
@@
TABLE V: Results on OGB link prediction datasets. Higher scores indicate better performance, with the best results highlighted in bold. Herein, ”HE” stands for Heuristic Encoding, ”X” represents node attributes, and ”NE” denotes Node Embedding.
@@
MF [menon2011link] $13.68\pm 4.75$ $38.86\pm 0.49$ $32.29\pm 0.94$ $51.86\pm 8.43$
FSSDNMF [chen2022link] $14.62\pm 2.64$ $37.95\pm 3.25$ $34.15\pm 1.16$ $54.71\pm 8.73$
DeepWalk [perozzi2014deepwalk] $26.42\pm 6.10$ $50.37\pm 0.34$ $28.88\pm 1.63$ $60.11\pm 0.23$
Node2vec [grover2016node2vec] $23.26\pm 2.09$ $48.88\pm 0.54$ $22.26\pm 0.83$ $61.41\pm 0.11$
NodePiece  [galkin2021nodepiece] $24.15\pm 3.04$ $47.88\pm 0.41$ $22.85\pm 0.94$ $61.52\pm 2.91$
GraphSAGE [sage2017inductive] $83.90\pm 4.74$ $48.10\pm 0.81$ $16.55\pm 2.40$ $82.60\pm 0.36$
GAT [gat2018graph] $95.38\pm 0.94$ $52.26\pm 0.85$ $51.33\pm 2.16$ $83.17\pm 0.54$
Neo-GNN [yun2021neo] $75.72\pm 3.42$ $55.31\pm 0.53$ $49.13\pm 0.60$ $87.26\pm 1.84$
PLNLP [wang2021pairwise] $90.88\pm 3.13$ $52.92\pm 0.98$ $32.38\pm 2.58$ $84.92\pm 0.29$
NBFNet [zhu2021neural] $18.14\pm 2.12$ $51.15\pm 1.38$ $23.96\pm 2.03$ $74.91\pm 2.37$
SEAL [zhang2018link] $30.56\pm 3.86$ $54.71\pm 0.79$ $48.80\pm 4.56$ $87.67\pm 0.32$
DEGNN [li2020distance] $26.63\pm 6.42$ $53.74\pm 0.45$ $36.48\pm 5.38$ $60.30\pm 0.81$
SIEG [ai2022structure] $31.95\pm 3.93$ $55.35\pm 0.52$ $53.35\pm 1.39$ $\mathbf{89.87\pm 0.10}$
@@
Case study 2. NBFNet lacks the algorithmic ability to leverage node embeddings. NBFNet [zhu2021neural] is a model specifically developed for link prediction. Differently from GNN-based link prediction methods like Algorithm 2 and SEAL-type methods, NBFNet generalizes traditional link heuristics such as Katz index [katz1953new], Personalized PageRank [langville2011google] into a general formulation and approximates such formulation using a special network. Unlike aggregation-based GNNs that propagate and aggregate node-wise representations, NBFNet is designed to train edge-wise representations. The model architecture of NBFNet makes it hardly consider node-wise information. This would make NBFNet lack the algorithmic ability to train powerful node embeddings and may lead to non-competitive link prediction performance on dense graphs, considering the strong performance of the GNN only using node embeddings on dense graphs as shown in Fig. 2.
VIII-C Further analysis of experimental results
@@
We expand our limitation analysis of existing link prediction methods by examining their experimental performance on four OGB benchmark datasets. The results are presented in Table V and Fig. 5. In the interest of brevity, our analysis focuses on several main types of link prediction methods.
@@
First, as shown in Fig. 5, the performance of heuristic methods is not stable across the four datasets. For example, RA performs best on ogbl-ppa but second worst on ogbl-ddi. These results are consistent with the research of [kovacs2019network] which indicates that many link heuristics are designed for specific applications and may perform well only on those specific graphs. Moreover, the unstable performance of every single heuristic confirms the need of combining multiple heuristics in link prediction, as shown in our Algorithm 2.
@@
We also report the heuristic encoding results (Fig. 6) obtained through Algorithm 2 when only employing heuristic encoding. We find that it is not always true that the more the heuristics used, the better the performance of HE. In contrast, encoding a certain number of heuristics can yield the best results, whereas encoding too many heuristics would degrade performance. This may be due to optimization challenges, such as certain heuristics introducing strong correlations or biases that lead the model to converge to suboptimal local minima.
@@
The node embedding methods based on relative distance encoding (DeepWalk [perozzi2014deepwalk], NodePiece [galkin2021nodepiece]) perform slightly better than those based on matrix factorization (MF [menon2011link], FSSDNMF [chen2022link]). Nevertheless, all these methods fall short compared to other methods. This could be attributed to the limitations of such methods, e.g., reliance solely on the adjacency matrix or unsupervised learning without link samples. This also underscores the critical role of link samples in supervising the training of node embeddings for link prediction.
@@
Fig. 5 also presents the results of MLP and general GNNs (GCN [kipf2017semi], GAT [gat2018graph] and JKNet [xu2018representation]) that use node embeddings only. MLP(NE) performs much worse than GNNs, demonstrating the significance of the neighborhood aggregation of GNNs in training node embeddings, considering that MLP updates each node’s representation independently of other nodes. Furthermore, GCN(NE) and GAT(NE) perform comparably, indicating that the expressiveness of GCN is sufficient for learning node embedding. The similar performance of GCN and GAT empirically supports our analyses in Remark 2 and 3, where we point out that the attention mechanism (e.g., GAT) cannot address the inherent issue of GNNs in learning structural information related to the number of each node’s neighbors and of common neighbors between two nodes.
@@
In Fig. 5, SEAL-type methods show state-of-the-art performance. Especially, SIEG [ai2022structure] achieves the best results on two sparse graphs, i.e., ogbl-collab and ogbl-citation2 with graph degrees of $8$ and $21$, respectively. However, they perform worse than general GNNs with node embeddings (GCN(NE) [kipf2017semi], GAT(NE) [gat2018graph] and JKNet(NE) [xu2018representation]) on two dense graphs, i.e., ogbl-ppa and ogbl-ddi. This discrepancy in the performance of SEAL-type methods could be attributed to the algorithmic challenge of training node embeddings using subgraphs. Unlike general GNNs, the algorithm of SEAL-type methods limits each node to perceive other nodes within the subgraph rather than the entire graph, thereby restricting the information flow between nodes and potentially reducing the efficiency of learning node embeddings.
@@
Besides, Fig. 5 shows two link prediction-specific methods, namely NBFNet [zhu2021neural] and Neo-GNN [yun2021neo]. NBFNet underperforms on four datasets, which aligns with the limitations identified in Section VIII-B. Neo-GNN predicts link likelihood by combining the scores obtained by heuristic methods and the result produced by a GNN. It performs on par with the state-of-the-art SEAL-type methods on two sparse graphs (ogbl-collab and ogbl-citation2).
@@
Lastly, our GNN(X)+HE based on Algorithm 2 performs better than SEAL-type methods on ogbl-collab, supporting our limitation analysis of SEAL-type methods in Section VIII-B, i.e., such methods could not effectively learn the information equivalent to NCN-dependent heuristics.
@@
IX Implication for Practical Applications